OpenStudy (anonymous):

Solve. 4x2 + 8x = 252 What are the solutions.

OpenStudy (anonymous):

X-7 & x+9

OpenStudy (phi):

I would divide all 3 terms by 4 to simplify the problem to \[ x^2 +2x= 63\] or, in standard form \[ x^2 +2x-63=0\] now list the factors of 63 1,63 3,21 7,9 the - sign on the last term means we are looking for a difference (the two factors have different signs) the + on the 2nd term means the larger factor is positive. 9 and 7 look good: choose 9 as + and 7 as - to get (x+9)(x-7)=0 now solve each in turn (x+9)=0 and (x-7)=0 can you list the 2 values of x that make these true?

OpenStudy (anonymous):

+1 medal for phi :-)